# Zero-Shot Multi-Object Shape Completion

### v1

@@ 5.1 Quantitative Results
To assess the importance of the large-scale 3D shape completion datasets, we train our model on splits of increasing sizes which contain $1$%, $5$%, $10$%, $20$%, $40$%, $60$%, $80$%, and $100$% of the total number of the objects in our dataset. We report metrics on the test split of our dataset. Figure 6 shows that all the metrics have a strong correlation with respect to the number of objects. This could imply that the model benefits significantly from increased data diversity and volume, enhancing its ability to understand and complete 3D shapes. We believe that this analysis is crucial for understanding the relationship between data quantity and model performance.
Figure 7: Qualitative results on our synthetic dataset (Top Left), YCB-Video (Top Right), HomebrewedDB (Bottom Left), and HOPE (Bottom Right) datasets. These results demonstrate the strong generalization to the real-world images on multi-object shape completion. We choose $3$ different views for better visibility.
@@ 5.2 Qualitative Results
Figure 7 shows the qualitative results of our method on both of the synthetic and real-world datasets from three different views. Unlike the synthetic dataset, the real-world depth measurements are more noisy and erroneous, however, we observe that our method can generate faithful and consistent 3D shapes on different types of objects. These results indicate that our model successfully learns geometric and semantic priors of real-world objects only from the synthetic data. Moreover, Figure 6 provides a comparison between our method and the second-best baseline, OCNN [63]. OCNN struggles with multi-object reasoning, resulting in unnatural artifacts improperly merging multiple objects. We believe this finding further supports that using 3D latent MAE helps capture a global context for better shape completion.
@@ 6 Conclusion and Future Work
In this paper, we present OctMAE, a hybrid architecture of Octree U-Net and latent 3D MAE, for efficient and generalizable shape completion. Further, we create the first large-scale and diverse 3D shape completion dataset, which consists of $1$M images rendered with $12$K objects with realistic scale. Our experimental results on a wide range of the datasets demonstrate accurate zero-shot multi-object shape completion is possible with a proper choice of the network architecture and dataset, which potentially facilitates several challenging robotics tasks such as robotic manipulation and motion planning. Although our method achieves superior performance, it comes with some limitations. First, truncated objects are not reconstructed properly since depth measurements are not available. We believe we can overcome this problem by incorporating techniques for query proposal [73] and amodal segmentation [79]. The second limitation is that the semantic information of completed shapes is not predicted. Although our focus in this work is geometric shape completion, we believe it is an interesting direction to integrate a technique from an open-vocabulary segmentation methods to obtain instance-level completed shapes. Third, our method does not handle uncertainty of surface prediction explicitly. In future work, we plan to extend our method to model uncertainty to improve the shape completion quality and diversity.
Figure 8: Comparisons on HomebrewedDB dataset (Top), and HOPE (Bottom) datasets. For better visibility, we show the generated and ground truth shapes. The top and bottom rows show an image from near camera and back views respectively. Compared to the other methods, our method predicts accurate and consistent shapes on a challenging shape completion task for novel objects.
@@
Zero-Shot Multi-Object Shape Completion
@@ 7 Implementation Details of Baselines
We use the implementation from https://github.com/NVlabs/VoxFormer. We make a single modification to adapt to multi-object shape completion. Unlike the original setting, a measured depth map is more accurate than an estimated one. Thus, we directly use the input depth map to extract query tokens in Stage $1$. In addition, we leverage trilinear interpolation to reconstruct a surface at LoD-$7$.

### v2

@@ 5.1 Quantitative Results
To assess the importance of the large-scale 3D scene completion datasets, we train our model on splits of increasing sizes which contain $1$%, $5$%, $10$%, $20$%, $40$%, $60$%, $80$%, and $100$% of the total number of the objects in our dataset. We report metrics on the test split of our dataset. Figure 6 shows that all the metrics have a strong correlation with respect to the number of objects. This could imply that the model benefits significantly from increased data diversity and volume, enhancing its ability to understand and complete 3D shapes. We believe that this analysis is crucial for understanding the relationship between data quantity and model performance.
Figure 7: Qualitative results on our synthetic dataset (Top Left), YCB-Video (Top Right), HomebrewedDB (Bottom Left), and HOPE (Bottom Right) datasets. These results demonstrate the strong generalization to the real-world images on multi-object scene completion. We choose $3$ different views for better visibility.
@@ 5.2 Qualitative Results
Figure 7 shows the qualitative results of our method on both of the synthetic and real-world datasets from three different views. Unlike the synthetic dataset, the real-world depth measurements are more noisy and erroneous, however, we observe that our method can generate faithful and consistent 3D shapes on different types of objects. These results indicate that our model successfully learns geometric and semantic priors of real-world objects only from the synthetic data. Moreover, Figure 6 provides a comparison between our method and the second-best baseline, OCNN [65]. OCNN struggles with multi-object reasoning, resulting in unnatural artifacts improperly merging multiple objects. We believe this finding further supports that using 3D latent MAE helps capture a global context for better scene completion.
@@ 6 Conclusion and Future Work
In this paper, we present OctMAE, a hybrid architecture combining an Octree U-Net and a latent 3D MAE, for efficient and generalizable scene completion. Further, we create the first large-scale and diverse 3D scene completion dataset, which consists of $1$M images rendered with $12$K objects with realistic scale. Our experimental results on a wide range of the datasets demonstrate accurate zero-shot multi-object scene completion is possible with a proper choice of the network architecture and dataset, which potentially facilitates several challenging robotics tasks such as robotic manipulation and motion planning. Although our method achieves superior performance, it comes with some limitations. First, truncated objects are not reconstructed properly since depth measurements are not available. We believe we can overcome this problem by incorporating techniques for query proposal [75] and amodal segmentation [81]. The second limitation is that the semantic information of completed shapes is not predicted. Although our focus in this work is geometric scene completion, we believe it is an interesting direction to integrate a technique from an open-vocabulary segmentation methods to obtain instance-level completed shapes. Third, our method does not handle uncertainty of surface prediction explicitly. In future work, we plan to extend our method to model uncertainty to improve the scene completion quality and diversity.
Figure 8: Comparisons on HomebrewedDB dataset (Top), and HOPE (Bottom) datasets. For better visibility, we show the generated and ground truth shapes. The top and bottom rows show an image from near camera and back views respectively. Compared to the other methods, our method predicts accurate and consistent shapes on a challenging scene completion task for novel objects.
Acknowledgment
We thank Zubair Irshad and Jenny Nan for valuable feedback and comments. This research is supported by Toyota Research Institute.
@@ Acknowledgment
Zero-Shot Multi-Object Scene Completion
@@ 7 Implementation Details of Baselines
We use the implementation from https://github.com/NVlabs/VoxFormer. We make a single modification to adapt to multi-object scene completion. Unlike the original setting, a measured depth map is more accurate than an estimated one. Thus, we directly use the input depth map to extract query tokens in Stage $1$. In addition, we leverage trilinear interpolation to reconstruct a surface at LoD-$7$.
@@ 9 Derivation of RoPE [62]
10 Additional Experiments
[…]
MCC [69] $1$ Ground truth $14.39$ $20.26$ $0.482$ $0.694$ $3.5\times 10^{4}$
ZeroShape [25] $1$ Ground truth $12.19$ $17.44$ $0.603$ $0.703$ $965.5$
Ours $1$ G-SAM [55] $14.98$ $20.73$ $0.666$ $0.683$ $240.3$
$N$ G-SAM [55] $13.56$ $16.19$ $0.700$ $0.699$ $83.4$
$N$ InSPyReNet [30] $12.31$ $14.47$ $0.724$ $0.712$ $84.0$
$1$ Ground truth $8.55$ $12.48$ $0.758$ $0.730$ $248.3$
$N$ Ground truth 6.97 9.31 0.803 0.750 85.1
Table 8: Ablations of mask sources and the number of target objects on the HOPE dataset.
10.1 Comparison against single-object methods
We trained our method, MCC [38] and ZeroShape [25] on our synthetic dataset with a single-object setup. For a fair comparison, we use ground-truth camera intrinsics and depth maps for Zeroshape. During evaluation, we complete each object individually and then concatenate all the completed objects in a scene. Table 8 demonstrates that our method with single- and multi-object setups outperforms the others regarding completion quality and runtime. Here, $1$ and $N$ in #Obj denote single- and multi-object setups, and $\text{CD}_{occ}$ is Chamfer distance of occluded surfaces. The large difference between CD and $\text{CD}_{occ}$ of single-object methods clearly show its poor occlusion handling due to the lack of multi-object reasoning.
10.2 Foreground vs Instance Masks
Zero-shot 2D instance segmentation of cluttered scenes is still challenging. For instance, the SoTA foreground detection model (InSPyReNet [30]) gives a $14.9\%$ higher IoU of a foreground mask than Grounded-SAM [55] (G-SAM), the latest zero-shot instance segmentation model, on HOPE dataset ($69.3\%$ vs $54.4\%$). For G-SAM, foreground masks are computed by combining its instance mask predictions, and its input prompt are manually tuned to improve an IoU. Further, Table 8 validates that using foreground masks during inference largely improves the final completion quality.